OpenStudy (anonymous):

help! find the limit. lim x--->0 1/x+4-14/and all over/x

OpenStudy (callisto):

Is this the question? \[\lim_{x \rightarrow 0} \frac{\frac{1}{x+4}-14}{x}\]

OpenStudy (anonymous):

yes but its -1/4 sorry :(

OpenStudy (callisto):

\[\lim_{x \rightarrow 0} \frac{\frac{1}{x+4}-\frac{1}{4}}{x}\]This?

OpenStudy (anonymous):

yes :)

OpenStudy (callisto):

\[\lim_{x \rightarrow 0} \frac{\frac{1}{x+4}-\frac{1}{4}}{x}\]\[=\lim_{x \rightarrow 0} \frac{\frac{(4)}{4(x+4)}-\frac{(x+4)}{4(x+4)}}{x}\]\[=\lim_{x \rightarrow 0} \frac{\frac{4-(x+4)}{4(x+4)}}{x}\]\[=\lim_{x \rightarrow 0} \frac{-x}{4x(x+4)}\]\[=\lim_{x \rightarrow 0} \frac{-1}{4(x+4)}\]Can you continue from here? Just put x=0 into the fraction..

OpenStudy (anonymous):

ok ummm yeah but will this give me -0.06?

OpenStudy (anonymous):

okay i have two more questions...do u mind?

OpenStudy (anonymous):

if i ask u 4 help?

OpenStudy (callisto):

It should be -1/16 .. :S Sorry.. I have something to do now... I'll be back later..

OpenStudy (anonymous):

okay thanks! :)

OpenStudy (callisto):

\[=\lim_{x \rightarrow 0} \frac{-1}{4(x+4)}\]\[= \frac{-1}{4(0+4)} \]\[= \frac{-1}{4(4)}\]\[= \frac{-1}{16}\]

OpenStudy (anonymous):

thank u! :)))